# PPT: Token Pruning and Pooling for Efficient Vision Transformers

### v1

@@ Appendix A DEEPER ANALYSIS
Figure 11: The average variance of the significance scores at each layer of different ViTs on the ImageNet validation set. Figure (a) and Figure (c) are generated based on the original models, while Figure (b) and Figure (d) are generated based on the compressed models with our method.
In Figure 11, we further refine our analysis by computing the average variance of significance scores across layers for different ViTs. We can observe a similar phenomenon across different ViTs, where the variance increases with layer depth, and compressed ViTs exhibit greater variance. This phenomenon to some extent validates the effectiveness and generality of our method.
Appendix B FULL RESULTS
Table 4: Results of PPT on different ViTs under various FLOPs, $*{}^{*}$ indicates that the ACC is evaluated under off-the-shelf (without fine-tuning).
@@ Appendix B FULL RESULTS
In addition, we provide the detailed data for plots in Figure 4 in Table 5 and Table 6.
@@
DynamicViT EViT ATS ToMe PPT (Ours)
2.0 - - 75.1 - 78.9
2.3 - 78.5 - - 79.2
2.5 - - 78.0 - 79.4
2.6 - 78.9 - 79.2 79.5
2.7 - - - 79.4 79.6
2.9 79.3 - - 79.5 79.8
3.0 - 79.5 79.7 - 79.8
3.4 79.6 - - 79.7 79.9
3.5 - 79.8 - - 79.9
4.0 79.8 79.8 - 79.7 80.0
@@ Appendix C EXTEND EXPERIMENTS
Different token pooling policy. Here we discuss different design choices for token pooling. For example, we could also utilize the efficient density peak clustering algorithm (DPC) (Rodriguez & Laio, 2014) for token merging. It identifies density peaks based on the local density and distance between image tokens, and assigns each token to the nearest density peak and forms clusters by merging nearby tokens that belong to the same peak. Figure 13 presents the effects of different token pooling policies. Our experiments demonstrate that BSM can improve the results of DPC by approximately 0.1% with fine-tuning, and the advantage gap is even larger without training.
@@
average of token similarity 2.9 79.5
variance of the significant score 79.2

### v2

@@ Appendix A Deeper Analysis
In Figure 11, we further demonstrate our analysis by computing the average variance of significance scores across layers for different ViTs. We can observe a similar phenomenon across different ViTs, where the variance increases with layer depth, and compressed ViTs exhibit greater variance. This phenomenon to some extent validates the effectiveness and generality of our method.
Figure 11: The average variance of the significance scores at each layer of different ViTs on the ImageNet validation set. Figure (a) and (c) are generated based on the original models, while Figure (b) and (d) are generated based on the compressed models with our method.
Table 4: Results of PPT on different ViTs under various FLOPs by adjusting the number of tokens removed in each stage, $*{}^{*}$ indicates that the ACC is evaluated under off-the-shelf (without fine-tuning).
@@
Appendix B Full Results
In addition, we provide the detailed data for plots in Figure 4, as shown in Table 5 and Table 6.
@@ Appendix B Full Results
Dynamic ViT
EViT ATS ToMe PPT (Ours)
2.0
75.1
78.9
2.3
78.5
79.2
2.5
78.0
79.4
2.6
78.9
79.2
79.5
2.7
79.4
79.6
2.9
79.3
79.5
79.8
3.0
79.5
79.7
79.8
3.4
79.6
79.7
79.9
3.5
79.8
79.9
4.0
79.8
79.8
79.7
80.0
@@ Appendix C Extend Experiments
Different token pooling policy. Here we discuss different design choices for token pooling. For example, we could also utilize the efficient density peak clustering algorithm (DPC) [22] for token merging. It identifies density peaks based on the local density and distance between image tokens, and assigns each token to the nearest density peak and forms clusters by merging nearby tokens that belong to the same peak. Figure 13 presents the effects of different token pooling policies. Our experiments demonstrate that BSM can surpass the results of DPC by approximately 0.1% with fine-tuning, and the advantage gap is even larger without training.
@@
average of token similarity 2.9 79.2
variance of the significant score 79.5